OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

2x+3y=5 so y=-(2/3)x+5/3 in the equation for a parallel line the slopes are the same, and we have the equation y=mx+b, m being the slope. -3=-(2/3)9+b -3=-6+b so b=3. the equation would then be y=-(2/3)x+3

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

you're welcome!